please help me find the solution of converting 365 days - seconds

Answer by Theo(13342)

365 * 24 = 8,760 hours
8760 * 60 = 525,600 minutes
525600 * 60 = 31,536,000 seconds
there are 24 hours in a day.
there are 60 minutes in an hour.
there are 60 seconds in a minute.